The average radius R of the orbit of a planet around the sun is determined by R = T2/3, where T is the number of years for one orbit and R is measured in astronomical units (AU). If it takes Pluto 248.53 years to make one orbit of the sun, then what is the average radius of the orbit of Pluto? If the average radius of the orbit of Neptune is 30.08 AU, then how many years does it take Neptune to complete one orbit of the sun?

Answer by Theo(3464)

If I understand the formula correctly, R = (2/3) * T

T is the number of years for one orbit.
R is the average radius of the orbit measured in Astronomical Units (AU).

Pluto takes 248.53 years to make one orbit of the sun.

What is the average radius of the orbit of Pluto.

Since R = (2/3) * T, then R = (2/3) * 248.53 = 165.6866667 AU.

If the average radius of the orbit of Neptune is 30.08 AU, then how many years does it take Neptune to complete 1 orbit of the sun?

Since R = (2/3) * T, and R = 30.08 AU, then the formula becomes:

30.08 = (2/3) * T

Solve for T to get T = 45.12 years